OpenStudy (anonymous):

Alfonso is 8 years older than Ashley. Five years ago Alfonso was three times as old as Ashley. How old is Ashley now 9 17 8

OpenStudy (anonymous):

98

OpenStudy (pfenn1):

Let x = Ashley's age now and y = Alfonso's age now. How old was Ashley 5 years ago? How old is Alfonso now? Can you write these expressions?

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

y=3x

OpenStudy (pfenn1):

So 5 years ago Ashley was x-5. What was Alfonso's age then? Note: Five years ago Alfonso was three times as old as Ashley.

OpenStudy (anonymous):

so =3x-5

OpenStudy (pfenn1):

Close. y-5 => Alfonso's age five years ago and 3 (x-5) => three times Ashley's age five years ago\[y-5=3(x-5)\]

OpenStudy (pfenn1):

What about Alfonso's age now in terms of Ashley's age now?

OpenStudy (anonymous):

y=3x-10

OpenStudy (anonymous):

right

OpenStudy (pfenn1):

That is the equation for theirs ages 5 years ago. However, now Alfonso is 8 years older than Ashley so \[y=8+x\] Understand?

OpenStudy (anonymous):

o_0

OpenStudy (pfenn1):

When you have two unknowns (x,y) you need two equations.\[y=3x-10\]\[y=8+x\] Now substitute the second equation into the first equation (or vice versa) for y.\[8+x=3x-10\]Now solve for x.

OpenStudy (anonymous):

thx x=9

OpenStudy (pfenn1):

Yep. So Alfonzo's age is 17 (Ashley's age plus 8) but 5 years ago, Ashley was 4 and Alfonso was 12, three times older than Ashley was then.

OpenStudy (pfenn1):

You're welcome.